Internal critical angle in core at core cladding interface of a step index fiber 78.6. The refractive index of cladding is 1.45 what is refractive index of core


1
Expert's answer
2022-06-13T08:24:00-0400
"\sin\theta_c=\frac{n_2}{n_1}"

"n_2=n_1\sin\theta_c\\
=1.45*\sin78.6^\circ=1.42"